Television sets: What does it mean to refer to a 20-in TV set or a 25-in TV set? A 15-in TV set also has a width of 12-inches. What is its height?
I am stumped on how to set up the problem into an equation. Can someone help me?

Answer by stanbon(75887)

What does it mean to refer to a 20-in TV set or a 25-in TV set?
It means the diagonal measurement is 20 in or 25 in.
---------------------------------------------------------
A 15-in TV set also has a width of 12-inches. What is its height?
Use Pythagoras to get:
height^2 + width^2 = diagonal^2
h^2 + 12^2 = 15^2
h^2 = 225-144
h^2 = 89
h = 9.43 inches
=====================
Cheers,
Stan H.